### v1

Abstract
Binary classification tasks with imbalanced classes pose significant challenges in machine learning. Traditional classifiers often struggle to accurately capture the characteristics of the minority class, resulting in biased models with subpar predictive performance. In this paper, we introduce a novel approach to tackle this issue by leveraging Universum points to support the minority class within quadratic twin support vector machine models. Unlike traditional classifiers, our models utilize quadratic surfaces instead of hyperplanes for binary classification, providing greater flexibility in modeling complex decision boundaries. By incorporating Universum points, our approach enhances classification accuracy and generalization performance on imbalanced datasets. We generated four artificial datasets to demonstrate the flexibility of the proposed methods. Additionally, we validated the effectiveness of our approach through empirical evaluations on benchmark datasets, showing superior performance compared to conventional classifiers and existing methods for imbalanced classification.
Keywords:
Binary classification, Quadratic support vector machine, Twin support vector machine, Class imbalance, Universum data.
1 Introduction
The support vector machine (SVM) model marked a pivotal advancement in binary classification, achieving this by identifying two parallel supporting hyperplanes that maximize the minimum margin between classes through a convex quadratic programming problem vapnik1974theory . Weston et al. introduced the Universum support vector machine ($\mathfrak{U}$-SVM), which enhances the model by incorporating prior information weston2006inference . This is achieved by including Universum samples that do not belong to any class. Additionally, the Twin Support Vector Machine (TSVM) represents a significant extension, utilizing two non-parallel hyperplanes to simultaneously separate the classes. This innovative approach maximizes the margin between them while ensuring sufficient coverage of both majority and minority class instances.
However, many real-world applications pose challenges as they often involve data that is not linearly separable. Addressing this requires robust techniques capable of handling nonlinear situations effectively. One well-established approach is to leverage the concept of kernel methods, which relies on the existence of a nonlinear mapping. This mapping transforms the original data into a higher-dimensional, potentially infinite-dimensional, feature space where the data becomes linearly separable. Notably, the appeal of kernel methods lies in their ability to bypass the need for explicitly knowing this mapping. Instead, they focus on selecting appropriate kernel functions with advantageous properties. However, the challenge lies in determining the suitable kernel function, and fine-tuning the associated hyperparameters can be computationally demanding. As a result, there is a growing interest in practical methods that seek nonlinear classifiers directly in the original feature space.
Kernel-free models offer a direct approach for handling nonlinearly separable datasets without the need to map them to a higher-dimensional feature space. The quadratic surface support vector machine (QSSVM) dagher2008quadratic utilizes quadratic surfaces for class separation. Bai et al.Bai2015 introduced a kernel-free least square QSSVM tailored for disease diagnosis. A kernel-free support vector machine, utilizing a double-well potential, proposed in gao2021kernel , aims to achieve a specialized fourth-order polynomial separating surface. However, when pursuing nonlinear classifiers, the complexity of the optimization program significantly increases due to the involvement of numerous decision variables, leading to computational challenges and potential overfitting. To mitigate this, incorporating a surrogate that promotes sparsity among decision variables proves advantageous moosaei2023sparse ; mousavi2019quadratic ; mousavi2020survey . To achieve less expensive quadratic programs, gao2019quadratic discusses a least squares twin QSSVM, leveraging two quadratic surfaces for data representation.
Class imbalance in binary classification presents numerous challenges that can significantly impact the performance and reliability of machine learning models haixiang2017learning . When one class dominates the dataset, classifiers tend to exhibit bias towards the majority class, resulting in subpar predictions for the minority class. This imbalance can lead to various issues, including reduced sensitivity and specificity, misinterpretation of model accuracy, and distorted decision boundaries. In critical domains such as fraud detection or medical diagnoses, where the minority class represents rare events, the consequences of misclassification can be severe fotouhi2019comprehensive ; makki2019experimental . Failing to address class imbalance may result in suboptimal model performance, compromised decision-making, and heightened risks of costly errors or missed opportunities. Thus, it is imperative to implement effective strategies to mitigate class imbalance and ensure that machine learning models accurately capture the underlying data patterns, irrespective of class distribution.
Our methodology draws inspiration from the approach introduced in richhariya2020reduced , which incorporates Universum points to bolster the minority class in linearly separable datasets. Building upon this foundation, we extend its application to the framework of quadratic twin support vector machines (QTSVM). Leveraging the inherent flexibility of QTSVM, which employs quadratic surfaces instead of hyperplanes for binary classification, we aim to enhance the efficacy of Universum points in addressing class imbalance. This innovative integration enables us to capture intricate class relationships and adapt to complex data distributions more effectively. In addition to the hinge loss, we derive the least square formulation, which adopts the $\ell_{2}$ norm to penalize slack variables and replaces inequality and equality constraints. To validate the effectiveness of our proposed methods, we conduct experiments on various datasets, including artificial and public benchmark datasets. These experiments demonstrate the efficacy and efficiency of our proposed models compared to other well-known SVM models.
Notation.
[…]
2 Problem Statement and Related Works
We start by describing the fundamental classical and kernel-free models proposed in the literature for binary classification when dealing with an (almost) linearly separable data set.
2.1 Support Vector Machine
Suppose that we are faced with a binary classification problem and the related data set is $T=\left\{(x_{1},y_{1}),\dots,(x_{m},y_{m})\right\}\in(\mathbb{R}^{n}\times\{\pm 1\})^{m},$ where the $x_{i}$s are $n$-dimensional samples and the $y_{i}$s are their corresponding labels. For such a two-class problem when the classes are (almost) linearly separable, the SVM algorithm finds two parallel hyperplanes with the maximal minimum margin. The training samples from the two classes are divided by the middle hyperplane $f(x)=w^{T}x+b=0$ if and only if $y_{i}(w^{T}x_{i}+b)\geq 1,\ i=1,\ldots,m.$ The hyperplane for the almost separable case can be obtained by solving the following optimization problem:
$\mathop{\min}\limits_{w,b,\xi}\,$ $\frac{1}{2}{{\left\|w\right\|}^{2}}+\mu\sum_{i=1}^{m}\xi_{i}$
s.t. $y_{i}(x^{T}_{i}w+b)\geq 1-\xi_{i},\quad i=1,\dots,m,$
$w\in\mathbb{R}^{n},\,b\in\mathbb{R},\ \xi\in\mathbb{R}^{m}_{+},$
where $\mu>0$ is a penalty parameter to control the trade-off and $\xi_{i}$’s are slack variables to accommodate noisy data points and outliers. The above soft margin SVM exhibits several desirable properties, including robustness to noise and outliers, as it minimizes the effect of individual misclassified points while maximizing the margin. However, addressing the dual of (2.1) is computationally advantageous due to its more structured nature. The Lagrangian dual problem is as follows:
$\mathop{\min_{\alpha}}$ $\frac{1}{2}\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{m}\alpha_{i}\alpha_{j}y_{i}y_{j}x_{i}^{T}x_{j}-\sum\limits_{i=1}^{m}\alpha_{i}$
s.t. $\sum\limits_{i=1}^{m}y_{i}\alpha_{i}=0,$
$0\leq\alpha_{i}\leq\mu,\quad i=1,\ldots,m,$
where $\alpha_{i}$ is the Lagrange multiplier. The decision function can be obtained by solving the above quadratic programming problem and is given by:
$f(x)=\operatorname{sgn}(w^{T}x+b)=\operatorname{sgn}\left(\sum_{i=1}^{m}\alpha_{i}y_{i}x_{i}^{T}x+b\right).$
A new sample is classified as $+1$ or $-1$ by utilizing the above decision function.
2.2 Twin Support Vector Machine
To optimize binary classification by identifying two hyperplanes that are maximally distant from each other and as close as possible to their respective classes, Jayadeva et al. khemchandani2007twin proposed the Twin Support Vector Machine (TSVM). This model delineates two non-parallel hyperplanes, as illustrated below:
$w_{1}^{T}x+b_{1}=0,\quad\mbox{and}\quad w_{2}^{T}x+b_{2}=0,$
where $w_{1},w_{2}\in\mathbb{R}^{n}$ and $b_{1},b_{2}\in\mathbb{R}$. Assume that the data points belonging to class $+1$ and class $-1$ are represented by the rows of matrices $A\in\mathbb{R}^{m_{1}\times n}$ and $B\in\mathbb{R}^{m_{2}\times n}$, respectively. Solving the following two quadratic programming problems yields the TSVM classifiers:
$\min_{w_{1},b_{1},\xi_{1}}$ $\frac{1}{2}\left\|Aw_{1}+e_{1}b_{1}\right\|^{2}+\frac{C_{1}}{2}e_{2}^{T}\xi_{1}$ (TSVM)
s.t. $-(Bw_{1}+e_{2}b_{1})+\xi_{1}\geq e_{2},\ \ \xi_{1}\geq 0,$
$\min_{w_{2},b_{2},\xi_{2}}$ $\frac{1}{2}\left\|Bw_{2}+e_{2}b_{2}\right\|^{2}+\frac{C_{2}}{2}e_{1}^{T}\xi_{2}$
s.t. $(Aw_{2}+e_{1}b_{2})+\xi_{2}\geq e_{1},\ \ \xi_{2}\geq 0,$
where $C_{1}$, $C_{2}>0$ are penalty parameters, $\xi_{1}$, $\xi_{2}$ are slack vectors, and $e_{1}$, $e_{2}$ are vectors of ones of appropriate dimension. Using the KKT conditions, the Wolfe dual formulations of problems in (TSVM) are
$\mathop{\max}_{\alpha}$ $\quad-\frac{1}{2}\alpha^{T}G(H^{T}H)^{-1}G^{T}\alpha+e_{2}^{T}\alpha$
s.t. $\quad 0\leq\alpha\leq C_{1}e_{2},$
$\mathop{\max}_{\alpha^{*}}$ $\quad-\frac{1}{2}\alpha^{*T}H(G^{T}G)^{-1}H^{T}\alpha^{*}+e_{1}^{T}\alpha^{*}$
s.t. $\quad 0\leq\alpha^{*}\leq C_{2}e_{1},$
where $H=[A~{}~{}e_{1}]$, $G=[B~{}~{}e_{2}]$ and $\alpha$, $\alpha^{*}$ are the Lagrangian multipliers. After solving the above dual problems, the separating hyperplanes are found by
$\begin{bmatrix}w_{1}\\ b_{1}\end{bmatrix}=-(H^{T}H)^{-1}G^{T}\alpha,\qquad\mbox{and}\qquad\begin{bmatrix}w_{2}\\ b_{2}\end{bmatrix}=(G^{T}G)^{-1}H^{T}\alpha^{*}.$
In practice, a small perturbation of the identity matrix is often added to ensure the invertibility of the coefficient matrices above. The following decision rule is used to assign a new data point $x\in\mathbb{R}^{n}$ to class $k\in\{1,2\}$:
$\mbox{class}~{}k=\mbox{argmin}\dfrac{|w_{i}^{T}x+b_{i}|}{\|w_{i}\|^{2}},~{}~{}~{}i=1,2.$
2.3 Universum Twin Support Vector Machine
Universum data is defined as a set of unlabeled samples that do not belong to any specific class sinz2007analysis ; weston2006inference . This type of data can encode past knowledge by providing meaningful information within the same domain as the problem at hand. Incorporating Universum data has been shown to effectively improve learning performance in both classification and clustering tasks. By integrating Universum data into classical models, researchers have enhanced the computational efficiency and generalization ability of various existing methods moosaei2023universum ; moosaei2023sparse ; weston2006inference ; xiao2020new .
Let $U\in\mathbb{R}^{r\times m}$ be a matrix where each row represents a Universum point. The Universum Twin Support Vector Machine ($\mathfrak{U}$-TSVM) was proposed to enhance the classification performance of (TSVM) qi2012twin . The $\mathfrak{U}$-TSVM was constructed by incorporating Universum data in the TSVM model, as the following pair of quadratic programming problems (QPPs):
$\mathop{\min}_{w_{1},b_{1},\xi_{1},\psi_{1}\,\,}$ $\frac{1}{2}{{\left\|Aw_{1}+e_{1}b_{1}\right\|}^{2}}+\dfrac{C_{1}}{2}e_{2}^{T}\xi_{1}+\frac{C_{u}}{2}e_{u}^{T}\psi_{1}$
s.t. $-(Bw_{1}+e_{2}b_{1})+{{\xi}_{1}}\geq e_{2},$
$(Uw_{1}+e_{u}b_{1})+\psi_{1}\geq(-1+\varepsilon)e_{u},$
$\xi_{1},\psi_{1}\geq 0,$
$\mathop{\min}_{w_{2},b_{2},\xi_{2},\psi_{2}\,\,}$ $\frac{1}{2}{{\left\|Bw_{2}+e_{2}b_{2}\right\|}^{2}}+\dfrac{C_{2}}{2}e_{1}^{T}\xi_{2}+\frac{C_{u}}{2}e_{u}^{T}\psi_{2}$
s.t. $(Aw_{2}+e_{1}b_{2})+{{\xi}_{2}}\geq e_{1},$
$-(Uw_{2}+e_{u}b_{2})+\psi_{2}\geq(-1+\varepsilon)e_{u},$
$\xi_{2},\psi_{2}\geq 0,$
where $C_{1},C_{2},$ and $C_{u}$ are positive penalty parameters, $\varepsilon\in(0,1)$ is the tolerance value for Universum class, $\xi_{1},\xi_{2},\psi_{1},$ and $\psi_{2}$ are measures of the violation of associated constraints. We may derive their dual problems by applying Lagrangian functions, as shown below.
$\mathop{\max}_{\alpha,\beta}$ $-\frac{1}{2}(G^{T}\alpha-O^{T}\beta)^{T}(H^{T}H)^{-1}(G^{T}\alpha-O^{T}\beta)+e_{2}^{T}\alpha+(\varepsilon-1)e_{u}^{T}\beta$
[…]
and
$\mathop{\max}_{\alpha^{*},\beta^{*}}$ $\ -\frac{1}{2}(H^{T}\alpha^{*}-O^{T}\beta^{*})^{T}(G^{T}G)^{-1}(H^{T}\alpha^{*}-O^{T}\beta^{*})+e_{1}^{T}\alpha^{*}+(\varepsilon-1)e_{u}^{T}\beta^{*}$
[…]
where $H=[A~{}~{}e_{1}],G=[B~{}~{}e_{2}]$, and $O=[U~{}~{}e_{u}]$. Once the above dual QPPs are solved, we can obtain the following parameters:
$\begin{bmatrix}w_{1}\\ b_{1}\end{bmatrix}=-(H^{T}H)^{-1}(G^{T}\alpha-O^{T}\beta),\quad\mbox{and}\quad\begin{bmatrix}w_{2}\\ b_{2}\end{bmatrix}=(G^{T}G)^{-1}(H^{T}\alpha^{*}-O^{T}\beta^{*}).$
Once these vectors are obtained, the separating planes $w_{1}^{T}x+b_{1}=0$ and $w_{2}^{T}x+b_{2}=0,$ are known. A new data point $x\in\mathbb{R}^{n}$ is allocated to class $i\in\{+1,-1\}$ by a similar rule to the TSVM.
2.4 Least-Square Twin Support Vector Machine
To improve the computational efficiency of a classifier, the Least-Square Twin Support Vector Machine (LS-TSVM) kumar2009least is introduced, inspired by TSVM. Unlike TSVM, LS-TSVM employs equality constraints rather than inequality constraints, leading to the solution of only a pair of linear equations. Typically, the linear LS-TSVM deals with the following pair of quadratic programming problems (QPPs)
$\mathop{\min}\limits_{w_{1},b_{1}\xi_{1}}$ $\frac{1}{2}{{\left\|Aw_{1}+e_{1}b_{1}\right\|}^{2}}+\dfrac{C_{1}}{2}\|\xi_{1}\|^{2}$ (LS-TSVM)
s.t. $-\left(Bw_{1}+e_{2}b_{1}\right)+\xi_{1}=e_{2},$
$\mathop{\min}\limits_{w_{2},b_{2},\xi_{2}}$ $\frac{1}{2}{{\left\|Bw_{2}+e_{2}b_{2}\right\|}^{2}}+\dfrac{C_{2}}{2}\|\xi_{2}\|^{2}$
s.t. $\left(Aw_{2}+e_{1}b_{2}\right)+\xi_{2}=e_{1},$
where $C_{1}$ and $C_{2}$ are positive penalty parameters. These convex programs have closed-form solutions.
Substituting the equality constraints of the above problems into associated objective functions, one obtains convex unconstrained optimization problems. Thus, to find their optimum solutions, we set their gradients with respect to $w_{1},b_{1},w_{2}$ and $b_{2}$ equal to 00. Consequently, we can find the two non-parallel hyperplanes by solving the following systems of linear equations
$\begin{bmatrix}w_{1}\\ b_{1}\end{bmatrix}=-(H^{T}H+C_{1}G^{T}G)^{-1}(C_{1}G^{T}e_{2}),$
$\begin{bmatrix}w_{2}\\ b_{2}\end{bmatrix}=(G^{T}G+C_{2}H^{T}H)^{-1}(C_{2}H^{T}e_{1}),$
using the same notation as before. A new data point $x\in\mathbb{R}^{n}$ is assigned to class one or two using a rule similar to that of the TSVM.
2.5 Least-Square Universum Twin Support Vector Machine
To enhance the generalization performance of LS-TSVM, Xu et al. xu2016least introduced the incorporation of Universum data, leading to the development of the Least-Square Universum Twin Support Vector Machine (LS-$\mathfrak{U}$-TSVM). This model is formulated as follows:
$\mathop{\min}_{w_{1},b_{1},\xi_{1},\psi_{1}\,\,}$ $\frac{1}{2}{{\left\|Aw_{1}+e_{1}b_{1}\right\|}^{2}}+\dfrac{C_{1}}{2}e_{2}^{T}\xi_{1}+\frac{C_{u}}{2}e_{u}^{T}\psi_{1}$ (LS-$\mathfrak{U}$-TSVM)
s.t. $-(Bw_{1}+e_{2}b_{1})+{{\xi}_{1}}=e_{2},$
$(Uw_{1}+e_{u}b_{1})+\psi_{1}=(-1+\varepsilon)e_{u},$
$\mathop{\min}_{w_{2},b_{2},\xi_{2},\psi_{2}\,\,}$ $\frac{1}{2}{{\left\|Bw_{2}+e_{2}b_{2}\right\|}^{2}}+\dfrac{C_{2}}{2}e_{1}^{T}\xi_{2}+\frac{C_{u}}{2}e_{u}^{T}\psi_{2}$
s.t. $(Aw_{2}+e_{1}b_{2})+{{\xi}_{2}}=e_{1},$
$-(Uw_{2}+e_{u}b_{2})+\psi_{2}=(-1+\varepsilon)e_{u},$
where $C_{1}$ and $C_{2}$ are positive penalty parameters, $\varepsilon\in(0,1)$ is the tolerance value for Universum class, $\xi_{1},\xi_{2},\psi_{1}$, and $\psi_{2}$ are measures of the violation of constraints associated. The unique solution to these problems is closed-form, which is advantageous for handling large-scale applications. By employing the same solution method as LS-TSVM, we can obtain the two non-parallel hyperplanes by solving the following system of linear equations.
$\begin{bmatrix}w_{1}\\ b_{1}\end{bmatrix}=-\big{(}H^{T}H+C_{1}G^{T}G+C_{u}$ $O^{T}O\big{)}^{-1}\big{(}C_{1}G^{T}e_{2}+C_{u}(1-\varepsilon)O^{T}e_{u}\big{)},$
$\begin{bmatrix}w_{2}\\ b_{2}\end{bmatrix}=\big{(}G^{T}G+C_{2}H^{T}H+C_{u}$ $O^{T}O\big{)}^{-1}\big{(}C_{2}H^{T}e_{1}+C_{u}(1-\varepsilon)O^{T}e_{u}\big{)}.$
Here, we use the same notation as for $\mathfrak{U}$-TSVM. A new data point $x\in\mathbb{R}^{n}$ is assigned to class $i\in\{+1,-1\}$ using a rule similar to that of the TSVM.
2.6 Quadratic Twin Support Vector Machine
Linear models, while effective for linearly separable data, may face limitations when dealing with more complex datasets. To address this challenge, kernel methods are introduced. However, these methods have their drawbacks gao2019quadratic . In response, many quadratic kernel-free methods have been recently studied mousavi2019quadratic ; gao2021novel ; gao2019rescaled . The idea behind these models is analogous to their linear version except they capture quadratic surfaces. Below, we elaborate on the Quadratic Kernel-Free Twin Support Vector Machine (QTSVM) model proposed by Gao et al. gao2019quadratic .
Let $\mathcal{I}_{1}:=\{i\in\{1,\dots,n\}\ |\ y_{i}=1\}$ and $\mathcal{I}_{2}:=\{i\in\{1,\dots,n\}\ |\ y_{i}=-1\}$. Suppose we have two classes that include $|\mathcal{I}_{1}|$ and $|\mathcal{I}_{2}|$ points, respectively. Then, the formulation of the QTSVM model for binary classification becomes the following:
$\min_{W_{1},b_{1},c_{1},\xi_{1}}$ $\sum_{i\in\mathcal{I}_{1}}(\frac{1}{2}x_{i}^{T}W_{1}x_{i}+b_{1}^{T}x_{i}+c_{1})^{2}+C_{1}\sum_{i\in\mathcal{I}_{2}}{\xi_{1}}_{i}$ (QTSVM)
subject to $1+(\frac{1}{2}{x_{i}^{T}W_{1}x_{i}+b_{1}^{T}x_{i}+c_{1})\leq{\xi_{1}}_{i},\quad\forall i\in\mathcal{I}_{2}}$
$W_{1}\in\mathbb{S}^{n},b_{1}\in\mathbb{R}^{n},c_{1}\in\mathbb{R},\xi_{1}\in\mathbb{R}^{|\mathcal{I}_{2}|}_{+}.$
$\min_{W_{2},b_{2},c_{2},\xi_{2}}$ $\sum_{i\in\mathcal{I}_{2}}(\frac{1}{2}x_{i}^{T}W_{2}x_{i}+b_{2}^{T}x_{i}+c_{2})^{2}+C_{2}\sum_{i\in\mathcal{I}_{1}}{\xi_{2}}_{i}$
subject to $1-(\frac{1}{2}{x_{i}^{T}W_{2}x_{i}+b_{2}^{T}x_{i}+c_{2})\leq{\xi_{2}}_{i},\quad\forall i\in\mathcal{I}_{1}}$
$W_{2}\in\mathbb{S}^{n},b_{2}\in\mathbb{R}^{n},c_{2}\in\mathbb{R},\xi_{2}\in\mathbb{R}^{|\mathcal{I}_{1}|}_{+}.$
where $C_{1}$ and $C_{2}$ are positive hyperparameters.
It is important to note that the quadratic-based models are not in the standard form of quadratic programs. Therefore, we will introduce some notations and provide several definitions to facilitate their conversion into standard quadratic program forms. We shall be succinct here as these definitions are mainly borrowed from mousavi2019quadratic . For a square matrix $A=[a_{ij}]_{i=1,\dots,n;j=1,\dots,n}\in\mathbb{R}^{n\times n}$, its vectorization given by
$\mbox{vec}(A):=\left[a_{11},\dots,a_{n1},a_{12},\dots,a_{n2},\dots,a_{1n},\dots,a_{nn}\right]^{T}\in\mathbb{R}^{n^{2}}.$
In case $A$ is symmetric, $\mbox{vec}(A)$ contains redundant information so that we often consider its half-vectorization given by:
$\mbox{hvec}(A):=\left[a_{11},\dots,a_{n1},a_{22},\dots,a_{n2},\dots,a_{nn}\right]^{T}\in\mathbb{R}^{\frac{n(n+1)}{2}}.$
Given $n\in\mathbb{N}$, there exist a unique elimination matrix $L_{n}\in\mathbb{R}^{\frac{n(n+1)}{2}\times n^{2}}$ such that magnus1980elimination
$L_{n}\mbox{vec}(A)=\mbox{hvec}(A);\quad\forall\,A\in\mathbb{S}^{n},$
and further, this elimination matrix $L_{n}$ has full row rank magnus1980elimination . Conversely, for any $n\in\mathbb{N}$, there is a unique duplication matrix $D_{n}\in\mathbb{R}^{n^{2}\times\frac{n(n+1)}{2}}$ such that
$D_{n}\mbox{hvec}(A)=\mbox{vec}(A);\quad\forall\,A\in\mathbb{S}^{n}\qquad\mbox{and}\qquad L_{n}D_{n}=I_{\frac{n(n+1)}{2}}.$
Definition 1.
For fixed $i\in[m]$ and $j\in[r]$, let
$s_{i}$ ::=:= $\frac{1}{2}\mbox{hvec}(x_{i}x_{i}^{T}),\qquad s_{j}:=\frac{1}{2}\mbox{hvec}(u_{j}u_{j}^{T}),\qquad r_{i}:=[s_{i};x_{i}],\quad r_{j}:=[s_{j};u_{j}],$
$w$ ::=:= $\mbox{hvec}(W),\quad z:=[w;b],\quad V:=\big{[}I_{\frac{n(n+1)}{2}}\,\,0_{\frac{n(n+1)}{2}\times n}\big{]},\qquad X_{i}:=I_{n}\otimes x_{i}^{T},$
$M_{i}$ ::=:= $X_{i}D_{n},\ \ H_{i}:=[M_{i}\,\,I_{n}],\,\ G:=2\sum_{i=1}^{m}H_{i}^{T}H_{i},\ \ X:=[x_{1}^{T};x_{2}^{T},\dots;x_{m}^{T};u_{1}^{T};u^{T}_{2};\dots;u_{r}^{T}].$
Consequently, for fixed $i=1,\dots,m$ and $j=1,\dots,r$, we get the following equations:
$\left\{\begin{array}[]{ll}\frac{1}{2}x_{i}^{T}Wx_{i}+x_{i}^{T}b+c=z^{T}r_{i}+c,&\\ \frac{1}{2}u_{j}^{T}Wu_{j}+u_{j}^{T}b+c=z^{T}r_{j}+c,&\\ Wx_{i}=X_{i}\mbox{vec}(W)\,=X_{i}D_{n}\mbox{hvec}(W)\,=\,M_{i}\mbox{hvec}(W)\,=\,M_{i}w,&\\ Wx_{i}+b=\,M_{i}w+I_{n}b\,=\,H_{i}z,&\\ \sum_{i=1}^{m}\|Wx_{i}+b\|^{2}=\sum_{i=1}^{m}(H_{i}z)^{T}(H_{i}z)\,=\,z^{T}(\sum_{i=1}^{m}(H_{i})^{T}H_{i})z=\frac{1}{2}z^{T}Gz.\end{array}\right.$
Note that subindices $i$ and $j$ are consistently used to correspond with data points and Universum points throughout this paper, respectively. Using the aforementioned notations, let $z_{1}$ and $z_{2}$ be the decision variables constructed using $(W_{1},b_{1})$ and $(W_{2},b_{2})$, respectively. These variables are related to the quadratic surfaces for class 1 and class 2, respectively, which will be captured by twin models discussed later in the paper. Consequently, we can express the equivalent problem of QTSVM in the standard form of a quadratic minimization problem as follows:
$\min_{z_{1},c_{1},\xi_{1}}$ $\sum_{i\in\mathcal{I}_{1}}(z_{1}^{T}r_{i}+c_{1})^{2}+C_{1}\sum_{i=1}^{|\mathcal{I}_{2}|}{\xi_{1}}_{i}$ (QTSVM’)
subject to $1+(z_{1}^{T}r_{i}+c_{1})\leq{\xi_{1}}_{i},\quad\forall i\in\mathcal{I}_{2}$
$z_{1}\in\mathbb{R}^{\frac{n(n+1)}{2}+n},c_{1}\in\mathbb{R},\xi_{1}\in\mathbb{R}^{|\mathcal{I}_{2}|}_{+}.$
$\min_{z_{2},c_{2},\xi_{2}}$ $\sum_{i\in\mathcal{I}_{2}}(z_{2}^{T}r_{i}+c_{2})^{2}+C_{2}\sum_{i=1}^{|\mathcal{I}_{1}|}{\xi_{2}}_{i}$
subject to $1-(z_{2}^{T}r_{i}+c_{2})\leq{\xi_{2}}_{i},\quad\forall i\in\mathcal{I}_{1}$
$z_{2}\in\mathbb{R}^{\frac{n(n+1)}{2}+n},c_{2}\in\mathbb{R},\xi_{2}\in\mathbb{R}^{|\mathcal{I}_{1}|}_{+}.$
After obtaining the solutions to these quadratic programs—either by solving them directly or through their dual formulations—the optimal solution to the problem (QTSVM) is denoted as $(z_{k}^{*},c_{k}^{*})=(W_{k}^{*},b_{k}^{*},c_{k}^{*})$ for $k=1,2$. A new data point $x$ is then assigned to class $k$ based on the following rule:
$\mbox{Class}\,k=\mbox{argmin}_{i=1,2}\frac{|z_{i}^{T}r(x)+c_{i}|}{\|H(x)z_{i}+c_{i}\|_{2}^{2}}=\mbox{argmin}_{i=1,2}\frac{|\frac{1}{2}x^{T}W_{i}x+b_{i}^{T}x+c_{i}|}{\|W_{i}x+b_{i}\|_{2}^{2}}.$
2.7 Least Square Quadratic Twin Support Vector Machine
Combining the ideas of (LS-TSVM) and (QTSVM), the following model was introduced in gao2021novel to capture two quadratic surfaces going through their corresponding classes yet distant from the other class:
$\min_{W_{1},b_{1},c_{1},\xi_{1}}$ $\sum_{i\in\mathcal{I}_{1}}(\frac{1}{2}x_{i}^{T}W_{1}x_{i}+b_{1}^{T}x_{i}+c_{1})^{2}+C_{1}\sum_{i=1}^{|\mathcal{I}_{2}|}{\xi_{1}}^{2}_{i}$ (LS-QTSVM)
subject to $1+(\frac{1}{2}{x_{i}^{T}W_{1}x_{i}+b_{1}^{T}x_{i}+c_{1})={\xi_{1}}_{i},\quad\forall i\in\mathcal{I}_{2}}$
$W_{1}\in\mathbb{S}^{n},b_{1}\in\mathbb{R}^{n},c_{1}\in\mathbb{R},\xi_{1}\in\mathbb{R}^{|\mathcal{I}_{2}|},$
$\min_{W_{2},b_{2},c_{2},\xi_{2}}$ $\sum_{i\in\mathcal{I}_{2}}(\frac{1}{2}x_{i}^{T}W_{2}x_{i}+b_{2}^{T}x_{i}+c_{2})^{2}+C_{2}\sum_{i=1}^{|\mathcal{I}_{1}|}{\xi_{2}}^{2}_{i}$
subject to $1-(\frac{1}{2}{x_{i}^{T}W_{2}x_{i}+b_{2}^{T}x_{i}+c_{2})={\xi_{2}}_{i},\quad\forall i\in\mathcal{I}_{1}}$
$W_{2}\in\mathbb{S}^{n},b_{2}\in\mathbb{R}^{n},c_{2}\in\mathbb{R},\xi_{2}\in\mathbb{R}^{|\mathcal{I}_{1}|}.$
This model efficiently separates nonlinear classes, offering both accuracy and speed. Its ability to handle complex relationships without significant computational overhead makes it versatile and suitable for real-world applications across various domains.
The standard quadratic form of the latter model is the following:
$\min_{z_{1},c_{1},\xi_{1}}$ $\sum_{i\in\mathcal{I}_{1}}(z_{1}^{T}r_{i}+c_{1})^{2}+C_{1}\sum_{i=1}^{|\mathcal{I}_{2}|}{\xi_{1}}^{2}_{i}$ (LS-QTSVM’)
subject to $1+(z_{1}^{T}r_{i}+c_{1})={\xi_{1}}_{i},\quad\forall i\in\mathcal{I}_{2}$
$z_{1}\in\mathbb{R}^{\frac{n(n+1)}{2}+n},c_{1}\in\mathbb{R},\xi_{1}\in\mathbb{R}^{|\mathcal{I}_{2}|},$
$\min_{z_{2},c_{2},\xi_{2}}$ $\sum_{i\in\mathcal{I}_{2}}(z_{2}^{T}r_{i}+c_{2})^{2}+C_{2}\sum_{i=1}^{|\mathcal{I}_{1}|}{\xi_{2}}^{2}_{i}$
subject to $1-(z_{2}^{T}r_{i}+c_{2})={\xi_{2}}_{i},\quad\forall i\in\mathcal{I}_{1}$
$z_{2}\in\mathbb{R}^{\frac{n(n+1)}{2}+n},c_{2}\in\mathbb{R},\xi_{2}\in\mathbb{R}^{|\mathcal{I}_{1}|}_{+}.$
These analogous models have closed-from solutions. For $k=1$ and $2$, let us introduce the following matrices:
$R_{k}=[r_{1},r_{2},\dots,r_{i},\dots,r_{|\mathcal{I}_{k}|}]\quad\textrm{and}\quad D_{k}=\operatorname{diag}(y_{1},y_{2},\dots,y_{|\mathcal{I}_{k}|}).$
Then, for $k=1$, solving the KKT conditions leads to:
$\begin{bmatrix}R_{1}R_{1}^{T}+C_{1}R_{2}R_{2}^{T}&R_{1}e_{1}+C_{1}R_{2}e_{2}\\ e_{1}^{T}R_{1}^{T}+C_{1}e_{2}R_{2}^{T}&|\mathcal{I}_{1}|+C_{1}|\mathcal{I}_{2}|\end{bmatrix}\begin{bmatrix}z_{1}\\ c_{1}\end{bmatrix}=\begin{bmatrix}C_{1}R_{2}D_{{2}}e_{2}\\ C_{1}D_{2}e_{2}\end{bmatrix}.$
And, for $k=2$, we get:
$\begin{bmatrix}R_{2}R_{2}^{T}+C_{2}R_{1}R_{1}^{T}&R_{2}e_{2}+C_{2}R_{1}e_{1}\\ e_{2}^{T}R_{2}^{T}+C_{2}e_{1}R_{1}^{T}&|\mathcal{I}_{2}|+C_{2}|\mathcal{I}_{1}|\end{bmatrix}\begin{bmatrix}z_{2}\\ c_{2}\end{bmatrix}=\begin{bmatrix}C_{2}R_{1}D_{{1}}e_{1}\\ C_{2}D_{1}e_{1}\end{bmatrix}.$
After solving these systems for both classes, given a new point $x$, the decision function is based on the same rule analogous to that of QTSVM.
2.8 $\mathfrak{U}$-QTSVM
To integrate prior information in the form of Universum points and effectively separate nonlinear datasets, we bring the following model:
$\min_{W_{1},b_{1},c_{1},\xi_{1}}$ $\sum_{i\in\mathcal{I}_{1}}(\frac{1}{2}x_{i}^{T}W_{1}x_{i}+b_{1}^{T}x_{i}+c_{1})^{2}+C_{1}\sum_{i=1}^{|\mathcal{I}_{2}|}{\xi_{1}}_{i}+{C_{u}}\sum_{j=1}^{r}{\psi_{1}}_{j}$
subject to $1+(\frac{1}{2}{x_{i}^{T}W_{1}x_{i}+b_{1}^{T}x_{i}+c_{1})\leq{\xi_{1}}_{i},\quad\forall i\in\mathcal{I}_{2}}$
$\frac{1}{2}u_{j}^{T}W_{1}u_{j}+b_{1}^{T}u_{j}+c_{1}\geq-1+\varepsilon-{\psi_{1}}_{j},\quad j=1,\dots,r,$
$W_{1}\in\mathbb{S}^{n},b_{1}\in\mathbb{R}^{n},c_{1}\in\mathbb{R},\xi_{1}\in\mathbb{R}^{|\mathcal{I}_{2}|}_{+},\psi_{1}\in\mathbb{R}^{r}_{+}.$
$\min_{W_{2},b_{2},c_{2},\xi_{2}}$ $\sum_{i\in\mathcal{I}_{2}}(\frac{1}{2}x_{i}^{T}W_{2}x_{i}+b_{2}^{T}x_{i}+c_{2})^{2}+C_{2}\sum_{i=1}^{|\mathcal{I}_{1}|}{\xi_{2}}_{i}+{C_{u}}\sum_{j=1}^{r}{\psi_{2}}_{j}$
subject to $1-(\frac{1}{2}{x_{i}^{T}W_{2}x_{i}+b_{2}^{T}x_{i}+c_{2})\leq{\xi_{2}}_{i},\quad\forall i\in\mathcal{I}_{1}}$
$-(\frac{1}{2}u_{j}^{T}W_{2}u_{j}+b_{2}^{T}u_{j}+c_{2})\geq-1+\varepsilon-{\psi_{2}}_{j},\quad j=1,\dots,r,$
$W_{2}\in\mathbb{S}^{n},b_{2}\in\mathbb{R}^{n},c_{2}\in\mathbb{R},\xi_{2}\in\mathbb{R}^{|\mathcal{I}_{1}|}_{+},\psi_{2}\in\mathbb{R}^{r}_{+}.$
In the standard quadratic form, the above model becomes:
$\min_{z_{1},c_{1},\xi_{1},\psi_{1}}$ $\sum_{i\in\mathcal{I}_{1}}(z_{1}^{T}r_{i}+c_{1})^{2}+C_{1}\sum_{i=1}^{|\mathcal{I}_{2}|}{\xi_{1}}_{i}+C_{u}\sum_{j=1}^{r}{\psi_{1}}_{j}$ ($\mathfrak{U}$-QTSVM’)
subject to $1+(z_{1}^{T}r_{i}+c_{1})\leq{\xi_{1}}_{i},\quad\forall i\in\mathcal{I}_{2}$
$z_{1}^{T}r_{j}+c_{1}\geq-1+\varepsilon-{\psi_{1}}_{j}$
$z_{1}\in\mathbb{R}^{\frac{n(n+1)}{2}+n},c_{1}\in\mathbb{R},\xi_{1}\in\mathbb{R}_{+}^{|\mathcal{I}_{2}|},\psi_{1}\in\mathbb{R}^{r}.$
$\min_{z_{2},c_{2},\xi_{2},\psi_{2}}$ $\sum_{i\in\mathcal{I}_{2}}(z_{2}^{T}r_{i}+c_{2})^{2}+C_{2}\sum_{i=1}^{|\mathcal{I}_{1}|}{\xi_{2}}_{i}+C_{u}\sum_{j=1}^{r}{\psi_{2}}_{j}$
subject to $1-(z_{2}^{T}r_{i}+c_{2})\leq{\xi_{2}}_{i},\quad\forall i\in\mathcal{I}_{1}$
$-(z_{2}^{T}r_{j}+c_{2})\geq-1+\varepsilon-{\psi_{2}}_{j}$
$z_{2}\in\mathbb{R}^{\frac{n(n+1)}{2}+n},c_{2}\in\mathbb{R},\xi_{2}\in\mathbb{R}^{|\mathcal{I}_{1}|}_{+},\psi_{2}\in\mathbb{R}^{r}.$
After solving these quadratic programs directly or through dual theory in convex optimization, a new point is assigned to a class following the same approach outlined for QTSVM.
2.9 LS-$\mathfrak{U}$-QTSVM
Incorporating Universum points may escalate the computational complexity of our models. Therefore, opting for a least squares version of the last model is often preferred. This model can be formulated as follows:
$\min_{W_{1},b_{1},c_{1},\xi_{1},\psi_{1}}$ $\sum_{i\in\mathcal{I}_{1}}(\frac{1}{2}x_{i}^{T}W_{1}x_{i}+b_{1}^{T}x_{i}+c_{1})^{2}+C_{1}\sum_{i=1}^{|\mathcal{I}_{2}|}{\xi_{1}}^{2}_{i}+{C_{u}}\sum_{j=1}^{r}{\psi_{1}}^{2}_{j}$ (LS-$\mathfrak{U}$-QTSVM)
subject to $1+(\frac{1}{2}{x_{i}^{T}W_{1}x_{i}+b_{1}^{T}x_{i}+c_{1})={\xi_{1}}_{i},\quad\forall i\in\mathcal{I}_{2}}$
$\frac{1}{2}u_{j}^{T}W_{1}u_{j}+b_{1}^{T}u_{j}+c_{1}=-1+\varepsilon-\psi_{j},\quad j=1,\dots,r,$
$W_{1}\in\mathbb{S}^{n},b_{1}\in\mathbb{R}^{n},c_{1}\in\mathbb{R},\xi_{1}\in\mathbb{R}^{|\mathcal{I}_{2}|},\psi_{1}\in\mathbb{R}^{r}.$
$\min_{W_{2},b_{2},c_{2},\xi_{2},\psi_{2}}$ $\sum_{i\in\mathcal{I}_{2}}(\frac{1}{2}x_{i}^{T}W_{2}x_{i}+b_{2}^{T}x_{i}+c_{2})^{2}+C_{2}\sum_{i=1}^{|\mathcal{I}_{1}|}{\xi_{2}}^{2}_{i}+{C_{u}}\sum_{j=1}^{r}{\psi_{2}}^{2}_{j}$
subject to $1-\big{(}\frac{1}{2}{x_{i}^{T}W_{2}x_{i}+b_{2}^{T}x_{i}+c_{2}\big{)}={\xi_{2}}_{i},\quad\forall i\in\mathcal{I}_{1}}$
$-\big{(}\frac{1}{2}u_{j}^{T}W_{2}u_{j}+b_{2}^{T}u_{j}+c_{2}\big{)}=-1+\varepsilon-\psi_{j},\quad j=1,\dots,r,$
$W_{2}\in\mathbb{S}^{n},b_{2}\in\mathbb{R}^{n},c_{2}\in\mathbb{R},\xi_{2}\in\mathbb{R}^{|\mathcal{I}_{1}|},\psi_{2}\in\mathbb{R}^{r}.$
The above model in the standard quadratic form is as follows:
$\min_{z_{1},c_{1},\xi_{1},\psi_{1}}$ $\sum_{i\in\mathcal{I}_{1}}(z_{1}^{T}r_{i}+c_{1})^{2}+C_{1}\sum_{i=1}^{|\mathcal{I}_{2}|}{\xi_{1}}^{2}_{i}+C_{u}\sum_{j=1}^{r}{\psi_{1}}^{2}_{j}$ (LS-$\mathfrak{U}$-QTSVM’)
subject to $1+(z_{1}^{T}r_{i}+c_{1})={\xi_{1}}_{i},\quad\forall i\in\mathcal{I}_{2}$
$z_{1}^{T}r_{j}+c_{1}=-1+\varepsilon-{\psi_{1}}_{j}$
$z_{1}\in\mathbb{R}^{\frac{n(n+1)}{2}+n},\,c_{1}\in\mathbb{R},\,\xi_{1}\in\mathbb{R}^{|\mathcal{I}_{2}|},\,\psi_{1}\in\mathbb{R}^{r}.$
$\min_{z_{2},c_{2},\xi_{2},\psi_{2}}$ $\sum_{i\in\mathcal{I}_{2}}(z_{2}^{T}r_{i}+c_{2})^{2}+C_{2}\sum_{i=1}^{|\mathcal{I}_{1}|}{\xi_{2}}^{2}_{i}+C_{u}\sum_{j=1}^{r}{\psi_{2}}^{2}_{j}$
subject to $1-(z_{2}^{T}r_{i}+c_{2})={\xi_{2}}_{i},\quad\forall i\in\mathcal{I}_{1}$
$-(z_{2}^{T}r_{j}+c_{2})=-1+\varepsilon-{\psi_{2}}_{j}$
$z_{2}\in\mathbb{R}^{\frac{n(n+1)}{2}+n},\,c_{2}\in\mathbb{R},\,\xi_{2}\in\mathbb{R}^{|\mathcal{I}_{1}|},\,\psi_{2}\in\mathbb{R}^{r}.$
The closed-form solutions to these problems are obtained as follows. For $k=1$ and $2$, we define:
$R_{k}=[r_{1},r_{2},\dots,r_{i},\dots,r_{|\mathcal{I}_{k}|}],\ \ D_{k}=\operatorname{diag}(y_{1},y_{2},\dots,y_{|\mathcal{I}_{k}|}).$
Also, let $U=[r_{1},r_{2},\dots,r_{j},\dots,r_{r}]$. Then, for $k=1$, the solution can be obtained by solving $A_{1}[z_{1};c_{1}]=b_{1}$ such that
$A_{1}=\begin{bmatrix}R_{1}R_{1}^{T}+C_{1}R_{2}R_{2}^{T}+C_{u}UU^{T}&R_{1}e_{1}+C_{1}R_{2}e_{2}+C_{u}Ue_{u}\\ e_{1}R_{1}^{T}+C_{1}e_{2}^{T}R_{2}^{T}+C_{u}e^{T}_{u}U^{T}&|\mathcal{I}_{1}|+C_{1}|\mathcal{I}_{2}|+rC_{u}\end{bmatrix}$
and
$b_{1}=\begin{bmatrix}C_{1}R_{2}D_{{2}}e_{2}+C_{u}(-1+\varepsilon)Ue_{u}\\ C_{1}D_{2}e_{2}+rC_{u}(-1+\varepsilon)\end{bmatrix},$
where $e_{u}$ is the all-one vector of the dimension of the Universum data. And, for $k=2$, the solution is obtained by solving $A_{2}[z_{2};c_{2}]=b_{2}$ where
$A_{2}=\begin{bmatrix}R_{2}R_{2}^{T}+C_{2}R_{1}R_{1}^{T}+C_{u}UU^{T}&R_{2}e_{2}+C_{2}R_{1}e_{1}+C_{u}Ue_{u}\\ e_{2}R_{2}^{T}+C_{2}e_{1}^{T}R_{1}^{T}+C_{u}e^{T}_{u}U^{T}&|\mathcal{I}_{2}|+C_{2}|\mathcal{I}_{1}|+rC_{u}\end{bmatrix}$
and
$b_{2}=\begin{bmatrix}C_{2}R_{1}D_{{1}}e_{1}-C_{u}(-1+\varepsilon)Ue_{u}\\ C_{2}D_{1}e_{1}-rC_{u}(-1+\varepsilon)\end{bmatrix}.$
After solving these quadratic programs, either directly or via dual theory in convex optimization, a new point is assigned to a class using the same approach described for QTSVM.
3 Imbalanced Universum Quadratic Twin Support Vector Machines
Class imbalance in binary classification poses significant challenges, leading to biased models and compromised predictive performance, particularly for minority classes. This imbalance can result in reduced sensitivity, misinterpretation of accuracy, and skewed decision boundaries, with severe consequences in critical domains like fraud detection or medical diagnoses. To address this, we draw inspiration from moosaei2023inverse ; richhariya2020reduced , incorporating Universum points to support minority classes in linearly separable datasets. Extending this to quadratic twin support vector machines, we leverage its flexibility to enhance Universum point effectiveness. Through this approach, we aim to improve classifier robustness and generalization, especially in imbalanced scenarios.
In the following parts, we maintain a general approach where the minority class ($A$) is considered as the positive class. To address the class imbalance, we employ random undersampling of the negative class ($B$) to create a balanced dataset in forming a quadratic surface for the minority class. Precisely, letting $A=\{x_{1},x_{2},\dots,x_{|\mathcal{I}_{1}|}\}$ and $B=\{\bar{x}_{1},\bar{x}_{2},\dots,\bar{x}_{|\mathcal{I}_{2}|}\}$ with $|\mathcal{I}_{1}|\ll|\mathcal{I}_{2}|$, we randomly select a reduced sample $\tilde{B}=\{\tilde{x}_{1},\tilde{x}_{2},\dots,\tilde{x}_{|\mathcal{I}_{1}|}\}$ from the negative class. We also construct $U=\{u_{1},u_{2},\dots,u_{r}\}$ where $r=|\mathcal{I}_{2}|-|\mathcal{I}_{1}|$ Universum points by averaging technique. Next, we choose a reduced sample of them and construct the reduced Universum sample $\hat{U}:=\{\hat{u}_{1},\hat{u}_{2},\dots,\hat{u}_{g}\}$ with $g=\lceil{|\mathcal{I}_{1}|}/{2}\rceil$. Note that $g\ll r$, and these values are deliberately chosen to ensure that the optimization problems associated with each class remain unbiased when an appropriate number of Universum points is added.
The main model of the paper presented next is designed to handle class imbalance when the positive class is assumed to be the minority. In this Imbalanced Universum Quadratic Twin Support Vector Machine (Im-$\mathfrak{U}$-QTSVM) model, we incorporate only $g$ Universum points in the formulation of the minority class to better refine the generalization boundaries in its favor. To address the bias caused by the majority class, we use $r$ Universum points in its formulation, almost as many as the number of points in this class. Slack variables are also used for penalizing misclassification in the class points and noise in the Universum points. Furthermore, an $\ell_{2}$ regularization term on the Hessian of the quadratic surface has been added to enhance the stability and generalization capability of the model. Consequently, the optimization problem for the Im-$\mathfrak{U}$-QTSVM model, which is based on the hinge loss function, can be formulated as follows:
$\min$ $\frac{1}{2}\sum_{i\in\mathcal{I}_{1}}(\frac{1}{2}{x_{i}^{T}W_{1}x_{i}+b_{1}^{T}x_{i}+{c_{1}}})^{2}+\frac{1}{2}C_{1}\sum_{i=1}^{|\mathcal{I}_{1}|}{\xi_{1}}_{i}+\frac{1}{2}C_{\hat{u}}\sum_{j=1}^{g}{\psi_{1}}_{j}+\frac{1}{2}\lambda_{1}\sum_{i\leq j}{W_{1}}^{2}_{ij}$
s.t. $-(\frac{1}{2}\tilde{x}_{i}^{T}W_{1}\tilde{x}_{i}+b_{1}^{T}\tilde{x}_{i}+c_{1})\geq 1-{\xi_{1}}_{i},\quad i=1,\dots,|\mathcal{I}_{1}|$
$(\frac{1}{2}\hat{u}_{j}^{T}W_{1}\hat{u}_{j}+b_{1}^{T}\hat{u}_{j}+c_{1})\geq-1+\varepsilon-{\psi_{1}}_{j},\quad j=1,\dots,g,$
$W_{1}\in\mathbb{S}^{n},\,b_{1}\in\mathbb{R}^{n},\,c_{1}\in\mathbb{R},\,\xi_{1}\in\mathbb{R}^{|\mathcal{I}_{1}|}_{+},\,\psi_{1}\in\mathbb{R}^{g}_{+}.$
$\min$ $\frac{1}{2}\sum_{i\in\mathcal{I}_{2}}(\frac{1}{2}{\bar{x}_{i}^{T}W_{2}\bar{x}_{i}+b_{2}^{T}\bar{x}_{i}+{c_{2}}})^{2}+\frac{1}{2}C_{2}\sum_{i=1}^{|\mathcal{I}_{2}|}{\xi_{2}}_{i}+\frac{1}{2}C_{u}\sum_{j=1}^{r}{\psi_{2}}_{j}+\frac{1}{2}\lambda_{2}\sum_{i\leq j}{W_{2}}^{2}_{ij}$
s.t. $\frac{1}{2}x_{i}^{T}W_{2}x_{i}+b_{2}^{T}x_{i}+c_{2}\geq 1-{\xi_{2}}_{i},\quad i=1,\dots,|\mathcal{I}_{1}|$
$\frac{1}{2}u_{j}^{T}W_{2}u_{j}+b_{2}^{T}u_{j}+c_{2}\geq 1-\varepsilon-{\psi_{2}}_{j},\quad j=1,\dots,r,$
$W_{2}\in\mathbb{S}^{n},\,b_{2}\in\mathbb{R}^{n},\,c_{2}\in\mathbb{R},\,\xi_{2}\in\mathbb{R}^{|\mathcal{I}_{1}|}_{+},\,\psi_{2}\in\mathbb{R}^{r}_{+}.$
Similarly, the least squares version of the model discussed above can manage class imbalance when the positive class is the minority. In the Imbalanced Least Square Universum Quadratic Twin Support Vector Machine (Im-LS-$\mathfrak{U}$-QTSVM) model, we also utilize $g$ Universum points in the minority class formulation to better refine the generalization boundaries in its favor. Similarly, to mitigate the bias caused by the majority class, we employ $r$ Universum points in its formulation, which is nearly equivalent to the number of points in this class. Slack variables are incorporated to penalize misclassification in the class points and noise in the Universum points as well. Additionally, an $\ell_{2}$ regularization term is added to the Hessian of the quadratic surface to improve the stability and generalization capability of the model. Further, we consider the quadratic function as the loss function. Therefore, the optimization problem for the Im-LS-$\mathfrak{U}$-QTSVM model can be expressed as follows:
$\min$ $\frac{1}{2}\sum_{i\in\mathcal{I}_{1}}(\frac{1}{2}{x_{i}^{T}W_{1}x_{i}+b_{1}^{T}x_{i}+{c_{1}}})^{2}+\frac{1}{2}C_{1}\sum_{i=1}^{|\mathcal{I}_{1}|}{\xi_{1}}^{2}_{i}+\frac{1}{2}C_{\hat{u}}\sum_{j=1}^{g}{\psi_{1}}^{2}_{j}+\frac{1}{2}\lambda_{1}\sum_{i\leq j}{W_{1}}^{2}_{ij}$
s.t. $1+\frac{1}{2}\tilde{x}_{i}^{T}W_{1}\tilde{x}_{i}+b_{1}^{T}\tilde{x}_{i}+c_{1}={\xi_{1}}_{i},\quad i=1,\dots,|\mathcal{I}_{1}|$
$-1+\varepsilon-(\frac{1}{2}\hat{u}_{j}^{T}W_{1}\hat{u}_{j}+b_{1}^{T}\hat{u}_{j}+c_{1})={\psi_{1}}_{j},\quad j=1,\dots,g,$
$W_{1}\in\mathbb{S}^{n},b_{1}\in\mathbb{R}^{n},c_{1}\in\mathbb{R},\xi_{1}\in\mathbb{R}^{|\mathcal{I}_{1}|},\psi_{1}\in\mathbb{R}^{g}.$
$\min$ $\frac{1}{2}\sum_{i\in\mathcal{I}_{2}}(\frac{1}{2}{\bar{x}_{i}^{T}W_{2}\bar{x}_{i}+b_{2}^{T}\bar{x}_{i}+{c_{2}}})^{2}+\frac{1}{2}C_{2}\sum_{j=1}^{|\mathcal{I}_{2}|}{\xi_{2}}^{2}_{i}+\frac{1}{2}C_{u}\sum_{j=1}^{r}{\psi_{2}}^{2}_{j}+\frac{1}{2}\lambda_{2}\sum_{i\leq j}{W_{2}}^{2}_{ij}$
s.t. $1-(\frac{1}{2}x_{i}^{T}W_{2}x_{i}+b_{2}^{T}x_{i}+c_{2})={\xi_{2}}_{i},\quad i=1,\dots,|\mathcal{I}_{1}|$
$1-\varepsilon-(\frac{1}{2}u_{j}^{T}W_{2}u_{j}+b_{2}^{T}u_{j}+c_{2})={\psi_{2}}_{j},\quad j=1,\dots,r,$
$W_{2}\in\mathbb{S}^{n},b_{2}\in\mathbb{R}^{n},c_{2}\in\mathbb{R},\xi_{2}\in\mathbb{R}^{|\mathcal{I}_{1}|},\psi_{2}\in\mathbb{R}^{r}.$
The constraints of the above least square version model are equations, so by substituting $\xi$ and $\psi$ into the objective function, and via Definition 1 and letting
$S_{A}$ $=[r_{1},r_{2},\dots,r_{|\mathcal{I}_{1}|}],$
$S_{B}$ $=[\bar{r}_{1},\bar{r}_{2},\dots,\bar{r}_{|\mathcal{I}_{2}|}],$
$S_{\tilde{B}}$ $=[\tilde{r}_{1},\tilde{r}_{2},\dots,\tilde{r}_{|\mathcal{I}_{2}|}],$
$S_{U}$ $=[r_{1},r_{2},\dots,r_{j},\dots,r_{r}],$
$S_{\hat{U}}$ $=[\hat{r}_{1},\hat{r}_{2},\dots,\hat{r}_{j},\dots,\hat{r}_{g}],$
the problem for the (first or) minority class can be reformulated as
$\min_{z_{1},c_{1}}$ $f=\dfrac{1}{2}\|S_{A}^{T}z_{1}+c_{1}e\|_{2}^{2}+\frac{C_{1}}{2}\big{\|}e+{S^{T}_{\tilde{B}}}z_{1}+c_{1}e\big{\|}^{2}_{2}$
$\qquad+\frac{C_{\hat{u}}}{2}\big{\|}(-1+\varepsilon)e-({S^{T}_{\hat{U}}}z_{1}+c_{1}e)\big{\|}^{2}_{2}+\frac{\lambda_{1}}{2}\|Vz_{1}\|^{2}_{2}.$ (Im-LS-$\mathfrak{U}$-QTSVM-1)
Here, $e$ is a vector of ones of appropriate size and $V$ was introduced in Def. 1. This problem can be solved by putting the gradient with respect to $z$ and $c$ equal to zero, so we have the following equations:
$\dfrac{\partial f}{\partial z}$ $=S_{A}(S_{A}^{T}z_{1}+c_{1}e)+C_{1}{S_{\tilde{B}}}(e+{S^{T}_{\tilde{B}}}z_{1}+c_{1}e)$ (1)
$~{}~{}~{}-C_{\hat{u}}{S^{T}_{\hat{U}}}((-1+\varepsilon)e-({S^{T}_{\hat{U}}}z_{1}+c_{1}e))+\lambda_{1}V^{T}Vz_{1}=0,$ (2)
$\dfrac{\partial f}{\partial c}$ $=e^{T}(S_{A}^{T}z_{1}+c_{1}e)+C_{1}e^{T}(e+{S^{T}_{\tilde{B}}}z_{1}+c_{1}e)-C_{\hat{u}}e^{T}((-1+\varepsilon)e-({S^{T}_{\hat{U}}}z_{1}+c_{1}e))=0.$ (3)
By integrating the above equations, we have the following system:
$\left[\begin{matrix}S_{A}S_{A}^{T}+C_{1}{S_{\tilde{B}}}{S^{T}_{\tilde{B}}}+C_{\hat{u}}{S_{\hat{U}}}{S^{T}_{\hat{U}}}+\lambda_{1}V^{T}V&Ae+C_{1}{S_{\tilde{B}}}e+C_{\hat{u}}{S_{\hat{U}}}e\\ e^{T}{S^{T}_{A}}+C_{1}e^{T}{S^{T}_{\tilde{B}}}+C_{\hat{u}}e^{T}{S^{T}_{\hat{U}}}&C_{1}e^{T}e+C_{\hat{u}}e^{T}e+|\mathcal{I}_{1}|\end{matrix}\right]\left[\begin{matrix}z_{1}\\ c_{1}\end{matrix}\right]\vspace{0.9cm}$
$=\left[\begin{matrix}-C_{1}{S_{\tilde{B}}}e+(-1+\varepsilon)C_{\hat{u}}{S_{\hat{U}}}e\\ -C_{1}e^{T}e+(-1+\varepsilon)C_{\hat{u}}e^{T}e\end{matrix}\right].$
Let
$\Sigma=\left[\begin{matrix}S_{A}S_{A}^{T}+C_{1}{S_{\tilde{B}}}{S^{T}_{\tilde{B}}}+C_{\hat{u}}{S_{\hat{U}}}{S^{T}_{\hat{U}}}+\lambda_{1}V^{T}V&Ae+C_{1}{S_{\tilde{B}}}e+C_{\hat{u}}{S_{\hat{U}}}e\\ e^{T}A^{T}+C_{1}e^{T}{S^{T}_{\tilde{B}}}+C_{\hat{u}}e^{T}{S^{T}_{\hat{U}}}&C_{1}e^{T}e+C_{\hat{u}}e^{T}e+|\mathcal{I}_{1}|\end{matrix}\right],$
and assume it is invertible. Let
$\beta=\left[\begin{matrix}-C_{1}{S_{\tilde{B}}}e+(-1+\varepsilon)C_{\hat{u}}{S_{\hat{u}}}e\\ -C_{1}e^{T}e+(-1+\varepsilon)C_{\hat{u}}e^{T}e\end{matrix}\right],$
then
$\left[\begin{matrix}z_{1}\\ c_{1}\end{matrix}\right]=\Sigma^{-1}\beta.$
Similarly, the problem corresponding to the majority class in the least square loss function becomes:
$\min_{z_{2},c_{2}}$ $\dfrac{1}{2}\|S^{T}_{B}z_{2}+c_{2}e\|_{2}^{2}+\frac{C_{2}}{2}\big{\|}e-(S_{A}^{T}z_{2}+c_{2}e)\big{\|}^{2}_{2}$
$+\frac{C_{u}}{2}\big{\|}(1-\varepsilon)e-(S_{U}^{T}z_{2}+c_{2}e)\big{\|}^{2}_{2}+\frac{\lambda_{2}}{2}\|Vz_{2}\|^{2}_{2},$ (Im-LS-$\mathfrak{U}$-QTSVM-2)
which by the same methodology leads to
$\tilde{\Sigma}=\left[\begin{matrix}S_{B}S^{T}_{B}+C_{2}S_{A}S^{T}_{A}+C_{u}{S_{U}S^{T}_{U}}+\lambda_{2}V^{T}V&S_{B}e+C_{2}S_{A}e+C_{u}S_{U}e\\ e^{T}S_{B}^{T}+C_{2}e^{T}S_{A}^{T}+C_{u}e^{T}S_{U}^{T}&C_{2}e^{T}e+C_{u}e^{T}e+|\mathcal{I}_{2}|\end{matrix}\right],$
and assume it is invertible. Let
$\tilde{\beta}=\left[\begin{matrix}C_{2}S_{B}S_{A}e+(1-\varepsilon)C_{u}S_{B}S_{U}e\\ C_{2}e^{T}e+(1-\varepsilon)C_{u}e^{T}e\end{matrix}\right],$
then
$\left[\begin{matrix}z_{2}\\ c_{2}\end{matrix}\right]=\tilde{\Sigma}^{-1}\tilde{\beta}.$
A new data point $x\in\mathbb{R}^{n}$ is assigned to class $i\in\{+1,-1\}$ using a rule similar to that of the QTSVM.
4 Theoretical Properties of Proposed Models
In this section, we present some theoretical properties of the proposed model. We first show that optimality is always ensured. Moreover, under very mild assumptions, the optimal solution is unique.
Theorem 4.1.
Problem (3) has always an optimal solution.
Proof.
The problem minimizes a convex quadratic function on an affine subspace, so it possesses an optimum. ∎
Theorem 4.2.
Problem (3) has a unique optimal solution if and only if the set of vectors $\{x_{i},\,i\in\mathcal{I}_{1};\;\tilde{x}_{i},\,i=1,\dots,|\mathcal{I}_{1}|;\;\hat{u}_{j},\,j=1,\dots,g\}$ is affinely independent.
Proof.
From Theorem 4.1, an optimum always exists. Recall that the problem has the form of minimization of a convex quadratic function on an affine subspace. Thus, the minimum is not unique if and only if the objective function is constant in the unbounded direction. This happens if and only if there is a nontrivial solution of
$\frac{1}{2}\tilde{x}_{i}^{T}W_{1}\tilde{x}_{i}+b_{1}^{T}\tilde{x}_{i}+c_{1}$ $={\xi_{1}}_{i},\quad i=1,\dots,|\mathcal{I}_{1}|$
$-(\frac{1}{2}\hat{u}_{j}^{T}W_{1}\hat{u}_{j}+b_{1}^{T}\hat{u}_{j}+c_{1})$ $={\psi_{1}}_{j},\quad j=1,\dots,g,$
for which the objective function vanishes. If the objective function is zero, then we can deduce that $\xi_{1}=0$, $\psi_{1}=0$ and $W_{1}=0$. Hence the remaining conditions take the form of
$b_{1}^{T}\tilde{x}_{i}+c_{1}$ $=0,\quad i=1,\dots,|\mathcal{I}_{1}|$
$b_{1}^{T}\hat{u}_{j}+c_{1}$ $=0,\quad j=1,\dots,g,$
$b_{1}^{T}{x}_{i}+c_{1}$ $=0,\quad\forall i\in\mathcal{I}_{1}.$
This characterizes linear dependence of vectors
$\{(x_{i};1),\,i\in\mathcal{I}_{1};\;(\tilde{x}_{i};1),\,i=1,\dots,|\mathcal{I}_{1}|;\;(\hat{u}_{j};1),\,j=1,\dots,g\},$
or, equivalently, affine dependence of the vectors in the formulation of the theorem. ∎
Naturally, the situation in which $z^{*}=0$ for an optimal solution $(z^{*},c^{*})$ is not desirable. Thus, we present a sufficient condition ensuring that the optimal solution is nonzero. This condition assumes that certain three vectors in space $\mathbb{R}^{\frac{n(n+1)}{2}}$ are linearly independent, which is very likely to be satisfied.
Proposition 1.
Let $(z^{*},c^{*})$ be an optimal solution of (3). If vectors $S_{A}e$, $S_{\tilde{B}}e$, and $S_{\hat{U}}e$ are linearly independent, then $z^{*}\not=0$.
Proof.
Suppose to the contrary that $z^{*}=0$. From equations (1) we obtain that
$c_{1}S_{A}e+C_{1}(1+c_{1}){S_{\tilde{B}}}e-C_{\hat{u}}(-1+\varepsilon-c_{1}){S^{T}_{\hat{U}}}e=0.$
This means that vectors $S_{A}e$, $S_{\tilde{B}}e$, and $S_{\hat{U}}e$ are linearly dependent; a contradiction. ∎
5 Numerical Experiments
In this section, a large scale of numerical experiments are conducted to validate the performance of the proposed Im-LS-$\mathfrak{U}$-QTSVM model for binary classification. We first introduce the experiment settings, including the data sources, the parameter setups, etc. Then all the numerical experiments are conducted to validate the classification accuracy of the proposed model along with some related binary classification models on some public benchmark data sets. Certain statistical tests are also conducted to analyze the computational results.
5.1 Experiment Settings
In the numerical experiments, multiple benchmark SVM models are implemented for comparison, including the linear SVM, linear Twin SVM and least square twin SVM models. We also tested their corresponding Universum variants. All the tested models and their abbreviations are listed in Table 1. In addition, we list the Python toolboxes (commercial solvers, packages, etc.) in Table 1. All the experiments are conducted on a personal computer with eight Intel (R) Core (TM) i7-13900 @ 3.4 GHz CPUs and 64GB RAM.
Table 1: Models implemented for the numerical experiments
Model Abbreviation Toolbox Parameters
SVM SVM Scikit-learn $C$
Twin SVM TSVM CVXOPT $C_{1},C_{2}$
Least Square Twin SVM LS-TSVM Numpy $C_{1},C_{2}$
Quadratic Least Square Twin SVM LS-QTSVM Numpy $C_{1},C_{2}$
Universum Twin SVM $\mathfrak{U}$-TSVM CVXOPT $C_{1},C_{2},C_{u},\varepsilon$
Least Square Universum Twin SVM LS-$\mathfrak{U}$-TSVM Numpy $C_{1},C_{2},C_{u},\varepsilon$
Universum Quadratic Twin SVM $\mathfrak{U}$-QTSVM CVXOPT $C_{1},C_{2},C_{u},\varepsilon$
Least Square Uni. Quadratic Twin SVM LS-$\mathfrak{U}$-QTSVM Numpy $C_{1},C_{2},C_{u},\varepsilon$
Proposed model Im-LS-$\mathfrak{U}$-QTSVM Numpy $C_{1},C_{2},C_{u},\lambda,\varepsilon$
For each experiment, a five-fold cross-validation procedure is applied. Each experiment is repeated ten times for each tested model in order to keep the results statistically meaningful. Means and standard deviations of the accuracy scores are recorded as the measurement of classification effectiveness. Ten percent of data samples from each class is randomly selected to generate Universum data by averaging each pair of samples from the two classes moosaei2023universum .
All the parameters for each model are tuned by applying the grid-search, which is commonly adopted for tuning SVM parameters gao2021novel ; luo2016soft . For each pair of the twin models in TSVM, LS-TSVM, LS-QTSVM, 2.3, LS-$\mathfrak{U}$-TSVM, 2.8, and LS-$\mathfrak{U}$-QTSVM, the penalty parameters $C_{1}$ and $C_{2}$ are selected to be the identical.
We first generate four artificial datasets to show the flexibility of the proposed Im-LS-$\mathfrak{U}$-QTSVM model. The surfaces produced by the proposed model are visualized in Figure 1, in which Figures 1(a) and 1(b) have the similar nonlinear pattern while Figures 1(c) and 1(d) have the other similar one. To show the proposed model on different levels of imbalanced data, we set the imbalanced rates in Figure 1(a) and 1(c) to be three while the imbalanced rates in Figure 1(b) and 1(d) to be ten. The major class of data is in color green while the minor class is in red. For the major class, the number of solid points is the same as the minor class.
(a) Arti-1, rate 1:3
(b) Arti-1, rate 1:10
(c) Arti-2, rate 1:3
(d) Arti-2, rate 1:10
Figure 1: Im-LS-U-QTSVM on artificial data sets
From the results in Figure 1 we notice that the proposed Im-LS-$\mathfrak{U}$-QTSVM model can fit the quadratic patterns. It also provides promising classification accuracy when the dataset is imbalanced under different levels.
5.2 Results on Public Benchmark Data
The proposed Im-LS-$\mathfrak{U}$-QTSVM model is tested on some public benchmark data sets. All the data information is listed in Table 2
Data set # of data points (Class 1/Class 2) # of features Imbalance rate
pima 768 (268/500) 8 1.87
blood transfusion 748 (178/570) 4 3.20
seismic bumps 2584 (170/2414) 15 14.20
haberman 306 (81/225) 3 2.78
wholesale 440 (142/298) 7 2.10
liver disorder 319 (134/185) 6 1.38
wine 130 (59/71) 13 1.20
balance 576 (288/288) 4 1.00
taxpayer 667 (331/336) 9 1.02
car evaluation 1594 (384/1210) 6 3.15
CTG 2126 (471/1655) 22 3.51
US heart 3658 (557/3101) 15 5.57
Alzheimer rowdel 354 (37/317) 9 8.57
Alzheimer clinical 1200 (373/827) 5 2.22
thyroid 215 (65/150) 5 2.31
ecoli 1 336 (143/193) 7 1.35
ecoli 123 336 (20/316) 7 4.25
ecoli 4 336 (64/272) 7 8.60
ecoli 1234 336 (9/327) 7 10.59
ecoli 5 336 (29/307) 7 15.80
ecoli 678 336 (35/301) 7 36.33
Table 2: Basic information of public benchmark data sets. [^2] The sources of data sets can be found here: https://github.com/tonygaobasketball/Sparse-UQSSVM-Models-for-Binary-Classification
Dataset Accuracy score (mean $\pm$ standard deviation) (%)
LS-$\mathfrak{U}$-QTSVM LS-$\mathfrak{U}$-TSVM LS-QTSVM LS-TSVM $\mathfrak{U}$-QTSVM $\mathfrak{U}$-TSVM TSVM SVM Im-LS-$\mathfrak{U}$-QTSVM
pima $77.74\pm 2.08$ $77.48\pm 1.97$ $77.48\pm 1.18$ $67.45\pm 0.93$ $77.48\pm 1.67$ $78.00\pm 2.10$ $70.19\pm 4.18$ $77.87\pm 1.63$ $\textbf{78.27}\pm 3.66$
blood transfusion $79.61\pm 1.61$ $79.48\pm 1.09$ $79.21\pm 1.91$ $76.61\pm 0.88$ $79.14\pm 1.39$ $78.34\pm 1.52$ $76.27\pm 0.27$ $76.27\pm 0.27$ $\textbf{80.42}\pm 3.14$
seismic bumps $93.32\pm 0.09$ $93.30\pm 0.18$ $92.86\pm 0.34$ $93.40\pm 0.16$ $93.11\pm 0.31$ $\textbf{93.42}\pm 0.06$ $90.71\pm 5.19$ $\textbf{93.42}\pm 0.01$ $\textbf{93.42}\pm 0.01$
haberman $76.96\pm 3.31$ $76.97\pm 2.56$ $75.49\pm 2.48$ $74.02\pm 0.77$ $75.49\pm 1.49$ $76.48\pm 4.05$ $73.69\pm 0.32$ $73.53\pm 0.48$ $\textbf{77.13}\pm 2.65$
wholesale $91.02\pm 2.87$ $91.93\pm 2.19$ $90.80\pm 2.28$ $69.20\pm 1.36$ $90.80\pm 2.12$ $91.70\pm 2.18$ $69.09\pm 1.26$ $91.59\pm 1.96$ $\textbf{92.05}\pm 2.25$
liver disorder $72.76\pm 5.61$ $69.61\pm 3.65$ $72.76\pm 5.45$ $59.87\pm 2.68$ $71.50\pm 5.32$ $70.39\pm 5.06$ $66.94\pm 4.71$ $69.45\pm 4.10$ $\textbf{73.69}\pm 4.69$
wine $78.85\pm 3.82$ $\textbf{100.00}\pm 0.00$ $76.54\pm 7.81$ $69.23\pm 6.81$ $95.38\pm 3.68$ $\textbf{100.00}\pm 0.00$ $70.38\pm 6.73$ $99.62\pm 0.77$ $\textbf{100.00}\pm 0.00$
balance $99.74\pm 0.52$ $95.66\pm 1.44$ $99.39\pm 0.49$ $65.01\pm 3.22$ $99.13\pm 0.93$ $95.75\pm 0.89$ $61.89\pm 2.51$ $95.48\pm 1.49$ $\textbf{99.78}\pm 0.22$
taxpayer $\textbf{55.40}\pm 3.56$ $53.90\pm 3.25$ $53.67\pm 1.67$ $50.37\pm 0.24$ $53.97\pm 3.23$ $53.82\pm 2.85$ $51.28\pm 2.44$ $51.28\pm 2.57$ $54.65\pm 3.28$
car evaluation $95.14\pm 0.73$ $86.54\pm 1.15$ $94.42\pm 0.93$ $80.40\pm 1.18$ $92.10\pm 0.94$ $86.54\pm 1.98$ $81.15\pm 1.54$ $86.20\pm 1.56$ $\textbf{95.26}\pm 0.92$
CTG $98.38\pm 0.55$ $96.05\pm 0.79$ $98.02\pm 0.62$ $79.42\pm 1.70$ $97.34\pm 0.63$ $96.73\pm 0.98$ $77.85\pm 0.07$ $97.06\pm 0.93$ $\textbf{98.42}\pm 0.63$
US heart $84.99\pm 0.80$ $\textbf{85.35}\pm 0.46$ $84.83\pm 0.31$ $84.94\pm 0.15$ $84.94\pm 0.21$ $85.27\pm 0.41$ $84.75\pm 0.18$ $84.77\pm 0.06$ $85.21\pm 0.48$
Alzheimer rowdel $90.40\pm 1.19$ $90.12\pm 1.72$ $89.55\pm 0.67$ $86.30\pm 2.59$ $88.71\pm 3.46$ $90.40\pm 1.68$ $75.40\pm 9.18$ $89.55\pm 0.67$ $\textbf{90.68}\pm 1.38$
Alzheimer clinical $77.00\pm 2.90$ $76.67\pm 2.28$ $75.92\pm 2.53$ $70.88\pm 1.23$ $76.54\pm 3.04$ $75.88\pm 2.32$ $70.42\pm 0.49$ $76.71\pm 2.29$ $\textbf{77.42}\pm 2.17$
thyroid $92.09\pm 5.22$ $89.77\pm 5.22$ $91.63\pm 4.56$ $75.81\pm 4.05$ $92.09\pm 4.31$ $91.16\pm 3.42$ $90.23\pm 5.58$ $88.84\pm 4.00$ $\textbf{95.35}\pm 2.55$
ecoli 1 $97.33\pm 1.73$ $97.33\pm 1.73$ $97.33\pm 1.73$ $67.55\pm 2.94$ $96.44\pm 2.41$ $97.03\pm 1.88$ $68.74\pm 1.09$ $96.73\pm 2.19$ $\textbf{97.92}\pm 1.52$
ecoli 123 $91.38\pm 1.92$ $90.20\pm 3.99$ $90.78\pm 1.69$ $80.95\pm 0.58$ $90.18\pm 2.57$ $89.60\pm 3.35$ $80.95\pm 0.58$ $89.29\pm 3.29$ $\textbf{91.97}\pm 2.20$
ecoli 4 $93.15\pm 1.22$ $93.15\pm 0.75$ $91.96\pm 1.81$ $81.27\pm 4.67$ $92.55\pm 1.65$ $92.84\pm 2.58$ $90.76\pm 2.92$ $92.85\pm 1.48$ $\textbf{94.04}\pm 1.35$
ecoli 1234 $97.91\pm 1.52$ $97.62\pm 1.52$ $97.91\pm 0.73$ $91.37\pm 0.59$ $96.73\pm 1.46$ $97.03\pm 1.62$ $91.37\pm 0.59$ $97.62\pm 1.20$ $\textbf{98.21}\pm 1.46$
ecoli 5 $\textbf{99.41}\pm 0.73$ $98.82\pm 2.35$ $98.81\pm 1.10$ $78.30\pm 5.17$ $98.51\pm 0.01$ $99.11\pm 1.18$ $94.05\pm 2.50$ $98.81\pm 1.10$ $\textbf{99.41}\pm 1.18$
ecoli 678 $99.11\pm 1.19$ $99.11\pm 0.96$ $98.51\pm 1.13$ $97.32\pm 0.59$ $98.81\pm 1.28$ $\textbf{99.41}\pm 0.96$ $97.32\pm 0.59$ $99.26\pm 0.96$ $\textbf{99.41}\pm 0.96$
Average rank of accuracy 2.71 3.95 4.76 8.00 5.19 4.05 8.19 5.48 1.14
Table 3: Public benchmark data accuracy results.
Dataset CPU time (s)
LS-$\mathfrak{U}$-QTSVM LS-$\mathfrak{U}$-TSVM LS-QTSVM LS-TSVM $\mathfrak{U}$-QTSVM $\mathfrak{U}$-TSVM TSVM SVM Im-LS-$\mathfrak{U}$-QTSVM
pima 0.141 0.013 0.128 0.181 0.537 0.671 0.798 0.035 0.199
blood transfusion 0.053 0.013 0.047 0.184 0.664 0.993 1.025 0.027 0.106
seismic bumps 1.690 0.079 1.555 4.766 41.329 33.805 54.145 0.126 2.847
haberman 0.019 0.008 0.016 0.043 0.097 0.152 0.178 0.009 0.032
wholesale 0.067 0.007 0.062 0.066 0.205 0.198 0.257 0.010 0.088
liver disorder 0.038 0.007 0.034 0.048 0.099 0.121 0.057 0.011 0.047
wine 0.065 0.004 0.065 0.008 0.082 0.021 0.019 0.004 0.070
balance 0.040 0.010 0.034 0.104 0.277 0.408 0.324 0.007 0.031
taxpayer 0.150 0.012 0.140 0.123 0.417 0.480 0.475 0.043 0.156
car evaluation 0.195 0.039 0.164 1.031 3.272 3.861 5.591 0.088 0.345
CTG 3.303 0.059 3.126 2.399 9.994 11.610 2.344 0.070 4.118
US heart 2.719 0.137 2.211 10.382 59.138 72.856 103.524 0.400 3.661
Alzheimer rowdel 0.081 0.007 0.077 0.056 0.230 0.252 0.243 0.011 0.141
Alzheimer clinical 0.114 0.024 0.096 0.590 1.965 2.174 2.189 0.070 0.178
thyroid 0.022 0.005 0.020 0.020 0.072 0.071 0.054 0.006 0.042
ecoli 1 0.052 0.007 0.050 0.045 0.140 0.130 0.119 0.005 0.061
ecoli 123 0.050 0.007 0.046 0.061 0.158 0.127 0.040 0.006 0.087
ecoli 4 0.051 0.007 0.053 0.035 0.176 0.274 0.219 0.005 0.099
ecoli 1234 0.052 0.006 0.046 0.025 0.186 0.155 0.062 0.005 0.096
ecoli 5 0.053 0.007 0.048 0.031 0.172 0.198 0.246 0.005 0.094
ecoli 678 0.051 0.007 0.049 0.049 0.239 0.205 0.054 0.004 0.097
Table 4: CPU time (s) consumed by implementing the model with fixed parameters.
For each tested model, the mean and the standard deviation of the classification accuracy on each data set are recorded in Table 3. The average rank of accuracy scores on all benchmark datasets is also calculated for each tested model. Moreover, the highest accuracy score on each benchmark dataset is highlighted. In addition, we have recorded the CPU time consumed by each tested model on all the benchmark datasets in Table 4.
Figure 2: Heatmap of relative rank of CPU time. For each row, the unit is closer to red if its CPU time is longer, while the unit is closer to green if the CPU time is shorter.
The observations from the results in Tables 3 and 4 can be summarized as follows.
  • Compared with all the other tested models, the proposed Im-LS-$\mathfrak{U}$-QTSVM achieves the highest mean accuracy scores on most of the benchmark datasets. Besides, it provides the highest average rank of accuracy score. In other words, the proposed model provides a better general classification accuracy than other tested models does in this experiment.
  • In general, the CPU time consumed by the Im-LS-$\mathfrak{U}$-QTSVM model is acceptable. As a least squares model, the proposed Im-LS-$\mathfrak{U}$-QTSVM does not lose efficiency when the size data set increases. Even for the relative large-scale datasets seismic bumps, US heart and CTG, the CPU time consumed by the proposed model is acceptable, and is less than that consumed by the other twin SVM models without least squares.
  • Furthermore, we plot a heatmap in Figure 2 to show the comparison of efficiency among all the tested models on each dataset. We may observe that, in most cases, the relative rank of the CPU time consumed by the proposed Im-LS-$\mathfrak{U}$-QTSVM model stays at the medium level among all the tested models. Compared with other tested models on a given dataset, the proposed model might be more efficient if the ratio $N/n$ is higher, where $n$ is the number of features and $N$ is the number of data samples.
5.3 Statistical analysis
Some statistical tests are conducted to further analyze the computational results we received in Section 5.2. We first conduct the Friedman test demvsar2006statistical to evaluate the classification accuracy. The null hypothesis is that all the tested models have the same classification accuracy.
Let $p$ and $q$ be the number of datasets and the number of tested models, respectively. Hence, $p=20$ and $q=9$, and the $\chi_{F}^{2}$ value of the Friedman test can be calculated by the formula (4)
$\chi_{F}^{2}=\frac{12p}{q(q+1)}\left(\sum_{i=1}^{q}R_{i}^{2}-\frac{q(q+1)^{2}}{4}\right)\approx 73.78,$ (4)
where $R_{i}$ is the average rank of accuracy by the $i$th model. And the statistics $F_{F}$ is then calculated in (5)
$F_{F}=\frac{(p-1)\chi_{F}^{2}}{p(q-1)-\chi_{F}^{2}}\approx 15.66,$ (5)
where $F_{F}$ is distributed according to the $F$-distribution with degrees of freedom $(8,160)$. With the significance level to be $\alpha=0.05$, the critical value of $F_{F}$ in our case is about 2.03, which is far less than the value of $F_{F}$. Therefore, we are confident in rejecting the null hypothesis. In other words, there exists a significant difference among all the tested models.
To further analyze the difference, the post-hoc Nemenyi test is performed as follows. The statistics CD is calculated by (6)
$CD=q_{.05}\sqrt{\frac{q(q+1)}{6p}}\approx 2.62,$ (6)
where $q_{.05}=3.10$ is the critical value of Tukey distribution. Therefore, the Nemenyi post-hoc test detects a significant difference between the proposed Im-LS-$\mathfrak{U}$-QTSVM model and each other tested model except the LS-$\mathfrak{U}$-QTSVM. The result from the Nemenyi post-hoc test is visualized in Figure 3.
Figure 3: Visualization of Nemenyi test.
Even though the average ranks of the proposed model and LS-$\mathfrak{U}$-QTSVM are close, we may still find the difference by the Wilcoxon signed-rank test as suggested in benavoli2016should . The null hypothesis is that there is no significant difference between the accuracy of the two models. Let $\beta_{j}$ be the difference between the accuracy of the two models on the $j$th dataset. The calculated values $R^{+}$ and $R^{-}$ defined as follows
$R^{+}$ $=\sum_{\beta_{i}>0}\text{rank}(\beta_{i})+\frac{1}{2}\sum_{\beta_{i}=0}\text{rank}(\beta_{i})=185,$
$R^{-}$ $=\sum_{\beta_{i}<0}\text{rank}(\beta_{i})+\frac{1}{2}\sum_{\beta_{i}=0}\text{rank}(\beta_{i})=25,$
Let $R_{\min}=\min\{R^{+},R^{-}\}=25$ and calculate the statistics $z$:
$z=\frac{R_{\min}-\frac{1}{4}p(p+1)}{\sqrt{\frac{1}{24}p(p+1)(2p+1)}}\approx-3.15<-1.96.$
Therefore, the null hypothesis can be rejected with a $0.05$ confidence level.
In conclusion, the proposed Im-LS-$\mathfrak{U}$-QTSVM model significantly outperforms all the other tested models with respect to classification accuracy on the benchmark datasets in Table 2.
6 Conclusions
In this paper, we introduced three kernel-free Universum quadratic surface support vector machine models for binary classification: $\mathfrak{U}$-QTSVM, LS-$\mathfrak{U}$-QTSVM, and Im-LS-$\mathfrak{U}$-QTSVM. Binary classification with imbalanced classes presents significant challenges in machine learning, especially when the minority class is underrepresented. Traditional classifiers often fail to accurately capture the characteristics of the minority class, leading to biased models with poor predictive performance. To address this issue, we developed the Im-LS-$\mathfrak{U}$-QTSVM model, tailored for imbalanced datasets. We examined several theoretical properties and proposed an efficient algorithm for this model. Our computational experiments confirmed the effectiveness and efficiency of the proposed models. The main findings are summarized below:
  • We incorporated Universum data into kernel-free QTSVM models to propose the $\mathfrak{U}$-QTSVM model, followed by its least-square version, LS-$\mathfrak{U}$-QTSVM.
  • To tackle the challenge of imbalanced datasets, we adapted $\mathfrak{U}$-QTSVM to create Im-$\mathfrak{U}$-QTSVM and its least-square version, Im-LS-$\mathfrak{U}$-QTSVM, which reduces computational complexity by solving a system of equations.
  • We explored several theoretical properties of the Im-LS-$\mathfrak{U}$-QTSVM model.
  • Numerical experiments demonstrated the impact of parameters on classification accuracy. The promising results on various artificial and public benchmark datasets indicate the effectiveness of the proposed models in addressing real-world binary classification problems.
Future research could extend this work in several ways. Immediate future work includes investigating the robustness of the proposed models for binary classification. Additionally, these models could be extended to semi-supervised learning scenarios and generalized for multi-class classification, which warrants further investigation.
Compliance with Ethical Standards
  • Conflict of interest: The authors declare that they have no conflicts of interest.
  • Ethical approval: This article does not contain any studies with human participants or animals performed by any of the authors.
  • Informed consents: Informed consent was obtained from all individual participants included in the study.
Authorship contributions: Hossein Moosaei: Conceptualization, Formal analysis, Investigation, Methodology, Project administration, Writing – review & editing. Milan Hladík: Conceptualization, Formal analysis, Project administration, Supervision, Validation, Writing – review & editing. Ahmad Mousavi: Formal analysis, Investigation, Writing, and Editing. Zheming Gao: Conceptualization, Data curation, Formal analysis, Investigation, Methodology, Software, Validation, Supervision, Writing – original draft. Haojie Fu: Software, validation, Writing – original draft.